the solution to this system of equations is (-2,1). Find the values for p and q.
px+(9-q)y=-10
(3p+1)x-(q-6)y=-21

Answer by jim_thompson5910(35256)

Hint:

Replace every copy of x with -2. Then replace every copy of y with 1. This will give you

p(-2)+(9-q)(1)=-10
(3p+1)(-2)-(q-6)(1)=-21

You now have a system of equations with 2 variables p and q. Solve this new system just like you would solve the old system with variables x and y.